OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 110r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of √r^5 square centimeters. Use the given information to determine an expression for the total number of pieces of paper used. Can someone please tel me how I would solve this. Or could you just write this in equation form, I have trouble doing so. Word problems just aren't my best. Please and thank you to whoever is able to help me! :)

OpenStudy (anonymous):

Does this sound good? Since each piece of paper has the same area ( √r^5 cm^2 ) You need to put √r^5 into exponential notation --> √r^5 = r^5/2 This can be thought of in two ways: a) the square root of r taken to the fifth power or b) finding the square root of the fifth power or r. they are equivalent Multiply it by a variable and set it equal to 100r^3 cm^2. N(r^(5/2) ) =100 r^3 N = 100 r63 / (r^(5/2)) = 100 r^(3 - (5/2)) = 100r^(1/2) N = 100 r^(1/2) = 100√r

OpenStudy (anonymous):

Yeah thats correct

OpenStudy (anonymous):

it sounds good...

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

TRUE STORY

OpenStudy (anonymous):

ur funny! :)